OpenStudy (anonymous):

In 1987 a recording artist sold 14,000 albums. In 1998 that same artist sold 4,683 albums. Find the rate of change in record sales from 1987 to 1998.

OpenStudy (anonymous):

so originally he sold 14000 is 1987 and 4683 in 1998. So change in sales = 14000 - 4683 = 9317 sales. rate of change = 9317 / 14000 x 100 %

OpenStudy (anonymous):

In 1987 a recording artist sold 14,000 albums. In 1998 that same artist sold 4,683 albums. Find the rate of change in record sales from 1987 to 1998. A.1/847 year per album years per album B. –847 albums per year C. 847 albums per year D. –847 years per album